Upd 218 controller
From DIDEAS Wiki
State / Motor Controller
- 1-3-09 : discovered jumper wire between TP94 and TP96 - this was shorting the AC enable to 3.3V. removed wire and now power board lites.
- 1-3-09 : this board is unable to generate an AC ENABLE to enable power board.
- 12-15-08 : force power on ECO & battery voltage ECO
- 12-14-08 : scan of SC ADCs show that they're out of range. (possibly this is a problem with TEST version of SC)
- 12-14-08 : applied power, MC is now operational
- 12-12-08 : MC switching regulator replaced???
- 12-11-08 : verified that MC switchers have Vsup. Applied SC 3.3V to MC 3.3V, ~80mA flowed, and the MC booted.
- Motor control NOT operational (State operational)
- replace U39
- 1-19-09 : believe to be installed into proto2 (since texas)
- 1-1-09 : believe to have problems with EEPROM???
- 12-10-08 : bench top offsets (3250, 3200) with power board #110-005
- 12-04-08 : Installed into PROTO2
- 12-04-08 : Board ran all night tests Good!
- 12-04-08 : started Upd_test_pattern_112608a with PB 110-008
- 12-04-08 : added force power on ECO and verified that power board lights
- 12-03-08 : U39 has been replaced with LT3470AEDDB#PBF
- 11-26-08 : NEEDS U39 replaced
- 11-25-08: Not operational
- 11-19-08 : traced MC power. Found that 3.3V supply only goes 1V - applied 3.3V from prg adpt and MC booted. Forced R155 high and 5V MC supply worked.
- 11-19-08 : appears that MC switcher is bad. Apply 3.3V to ISP, SC boots, but MC does not.
- 11-8-08 : requires external device to hold it on...
- 10-6-08 : installed in UP#3
- 1-8-09 : proto1 failed, although this board is 'operational' - may have problems that result in instability that cause the failures.
- 1-6-09 : used for motor controller dev CG6, many hours of successful benching testing. One problem is that after reprogramming the motor controller, there is a ADC offset on the MC. This is fixed by a power cycle.
- 1-3-09 : seems to power on - need to test
- 1-3-09 : lots of rework around SC turn on circuit. Thinking that Q5 is bad since R141 is grounded, R140 has Vin, but TP93 is jumping around as the system powers on. Will add a jumper wire from Vin to C104
- 1-3-09 : replaced broken wire for battery voltage ECO
- 12-17-08 : has trouble starting.
- 12-15-08 : Installed R192 replace C137, C146 with 0.0001uf
- 12-11-08 : current offsets caused by missing R192. Also verify filter caps - C137, C146 (remove or replace)
- 12-10-08 : offsets (3254, 22197) mode 7adc reads (2252, 3440)
- 11-30-08 : 1st current sensor has large offset
- 11-30-08 : appears that solder on a 1210 cap was making case contact. This was corrected, and tape is being put on the case surface.
- 11-28-08 : installed into #2, appears to have trouble turning on - staying on. Appears to be caused by rework and raised components shorting to case.
- 11-28-08 : Ran 36 hrs test with Upd_pcb#110-005
- 11-26-08 : operational
- 11-26-08 : R192 missing this is a pull up on the real time clock
- 11-26-08 : added R141 to GND ECO (force power on)
- 11-25-08 : Applied power, not operational
- 11-19-08 : Applied power, seems to run okay.
- 11-19-08 : Needs 3.3V to poweron status : component gets hot.
- 10-8-08 : received from UPD - tested - seems to be DOA
- 10-7-08 : received by UPD
- 10-6-08 : died in system while using the battery and under high current load - traced to 3.3V motor controller supply not working. Was used with power board that has TVS installed on caps, and external TVS were also used. Died
- 10-5-08 : installed in UP#3, made battery measurement ECO
- jumper wire between TP84 and high side of C104 (or C105)
- has battery voltage ECO
- has force power on ECO
- 1-19-09 : running great inside proto1
- 1-11-09 : ran overnight driving powerboard 110-08. Will install into proto1
- 1-9-09 : Replaced U39 and tests that SC LED flashes quickly and the MC led flashes slowly
- 1-3-09 : U39 needs to be replaced -
- 12-14-08 : powered and no motor controller LED
- 12-12-08 : switching regulator replaced??
- 12-11-08 : verified that MC switchers have Vsup. Applied SC 3.3V to MC 3.3V, ~100mA flowed, and the MC booted.
- 12-10-08 : MC is unpowered. Need to replace MC 3.3V switcher?
- 11-26-08 : appears that the motor controller dsPIC is unpowered.
- 11-19-08 : Rob added GND connection to keep board powered on (either at TP93 or otherside R141)
- 11-19-08 : Applied 3.3V to ISP header and SC booted.
- 11-19-08 : IMU #2 connected, push switch and won't power up
- 10-5-08 : need to evaluate
- 10-4-08 : Removed from UP#3, as randomly powers on at ~10hz due to noise? possibly damaged due to the pushbutton that WAS attached.
- 10-1-08 : Modified for auto-off : requires external signal for power on, has MC battery voltage ECO
- Test Run Board
==110-009 == (BF1)
- 12-22-08 : Replaced cap C142 & ECO battery voltage done
- 12-12-08 : robert states that there is mechanical damage to the PCB and some of the encoder decoders.
- 12-12-08 : still running in morning
- 12-11-08 : ran Upd_test_pattern_112608a with PB 110-008 overnight
- 12-11-08 : Added jumper wire from R68 to U16.6. Encoder now works.
- 12-11-08 : No QEI signal MC pic pins 11 and 13. Look back at U16,17. No signal on U16.6
- 12-10-08 : appears that board is unable to read what appears to be a working encoder.
- 12-10-08 : With FET 110-008, mode 7 offset (3128,3153) adc (2244, 2252)
- 12-04-08 : was operational in PROTO2 since Monday. Worked well. Removed to add ECO
- 11-29-08 : operational - tested with PB 100-002
- 11-26-08 : operational, but missing CAP in middle row (C142?)
11-20-08 : Needs the 24 pin flex connect to be replaced. (replaced 11-25-08)
- 11-19-08 : Needs battery measurement modification.
- 11-8-08 : has jumper wire to turn self on, but wouldn't boot until IMU attached. After first power, remove IMU and works fine.
- 1-21-09 : based overnight - installed into 'protoBAK'
- 1-20-09 : run over night test with PB-110-010
- 1-8-09 : believe is installed in proto1
- 12-19-08 : running over night Upd_test_pattern_112608a w/ PB 110-008
- 12-15-08 : completed over night Upd_test_pattern_112608a w/ PB 110-008
- 12-14-08 : removed from proto3. running test program
- 12-11-08 : installed in left (#3) - mode 7 offset without motor 3251 3235 (FET 110-004)
- 11-30-08 : verified with Upd_test_pattern_112608a for 3 hrs, using PB 110-004
- 11-30-08 : CEB added jumper to keep power on.
- 11-29-08 : seems operational
- 11-26-08 : needs 3.3V to be applied to get started.
- 11-25-08 : flex connector replaced
- replace ENC connector (bent pins), remove HALL connector